OpenStudy (anonymous):

Can anyone help me find the center, vertices, and foci of the ellipse? 4. x^2+4y^2+8x-48=0

terenzreignz (terenzreignz):

Well of course. ^_^ Now, it might help if we have a goal in mind, and that goal is to play with this equation so that it looks something like this: \[\Large \frac{(x-\color{red}h)^2}{\color{blue}a^2}+\frac{(y-\color{green}k)^2}{\color{orange}b^2}= 1\] You follow?

OpenStudy (anonymous):

yes.

terenzreignz (terenzreignz):

Okay, first, bring everything with no variable next to it to the right-side. Can you do that?

OpenStudy (anonymous):

yes .

terenzreignz (terenzreignz):

Then please do it, and tell me what you get.

OpenStudy (anonymous):

x^2+8x+4y^2=48

terenzreignz (terenzreignz):

Okay good :) \[\Large x^2 + 8x + 4y^2 = 48\] Now, the y-part has nothing to worry about for now, what about the x-part? You need to do what's called "completing the square" heard of it?

OpenStudy (anonymous):

yes.

terenzreignz (terenzreignz):

Then can you complete the square here on this part \[\Large \color{blue}{x^2 + 8x} + 4y^2 = 48\]

OpenStudy (anonymous):

there is nothing to complete because the x value is already 1

terenzreignz (terenzreignz):

No, I mean, for example, \[\Large z^2 + 6z\] to complete the square, you just have to take the coefficient of the variable without a square (the 6z) so that coefficient is 6. Get half of it, and then square it. Half of 6 is 3, and the square of 3 is 9. Now add and subtract that to the expression \[\Large z^2 + 6z \color{red}{+9 - 9}\] and you'll notice that this part \[\Large \color{blue}{z^2 +6z +9 }-9\] is now a perfect square:\[\Large \color{blue}{(z+3)^2}-9\]

terenzreignz (terenzreignz):

\[\Large \color{blue}{x^2 + 8x} + 4y^2 = 48\] do the same here.

OpenStudy (anonymous):

i guess x^2+8x+16-16 = (x+4)-16

terenzreignz (terenzreignz):

That's correct, except you're missing an exponent :P No matter...\[\Large \color{blue}{(x+4)^2-16} + 4y^2 = 48\]

terenzreignz (terenzreignz):

Now, there's that -16, mind bringing that to the right-side too? It has no variable after all...

OpenStudy (anonymous):

(x+4)^2+4y=64

terenzreignz (terenzreignz):

This time your y is missing an exponent :P But again, no matter, \[\Large (x+4)^2 + 4y^2 = \color{blue}{64}\]

terenzreignz (terenzreignz):

Ready for the next step? Remember that we want it to look like this: \[\Large \frac{(x-\color{red}h)^2}{\color{blue}a^2}+\frac{(y-\color{green}k)^2}{\color{orange}b^2}= 1\]so we want the right-side to be just 1. How do we do that?

OpenStudy (anonymous):

I guess by dividing it all by 64 ? ?

terenzreignz (terenzreignz):

Yup.\[\Large \frac{(x+4)^2}{64}+ \frac{4y^2}{64}=1\] simplify?

OpenStudy (anonymous):

(x+4)^2/64+y^2/16=1 ? correct ?

terenzreignz (terenzreignz):

Yup. Now all you have to do is identify your h,k, a and b.

terenzreignz (terenzreignz):

Here, a reference: \[\Large \frac{(x-\color{red}h)^2}{\color{blue}a^2}+\frac{(y-\color{green}k)^2}{\color{orange}b^2}= 1\]

OpenStudy (anonymous):

center is (-4,0) a=8 b=4 the major axis is x the vertices is (4,0) && (-12,0) && the foci is (-4,0 +(-)4sqrt3 . correct ?

terenzreignz (terenzreignz):

Impressive :D

OpenStudy (anonymous):

can you help me with (x+6)^2/12+(y-4)^2/16=1 -- i cant figure out the 12.

terenzreignz (terenzreignz):

sorry about that, numbers aren't always going to be pretty... you'll just have to make do with the fact that \(\large a = \sqrt{12}\) no other way, sorry :)

OpenStudy (anonymous):

ok , thanks !

terenzreignz (terenzreignz):

no problem :)